# On the Linear Speedup of the Push-Pull Method for Decentralized Optimization over Digraphs

### v1

@@ 1.1 Decentralized algorithms over digraphs
The second family consists of algorithms that use only row-stochastic matrices for communication, referred to as the Pull-Only setting. Just as Push-Sum underpins the design of Col-Only algorithms, the Pull-Diag gossip protocol [9] serves as the foundation for Row-Only methods. Building on Pull-Diag, reference [9] adapted the distributed gradient descent algorithm to the Row-Only setting. Subsequently, [18, 19] extended gradient tracking techniques to this framework, while momentum-based Pull-Only gradient tracking methods were developed by [20, 21].
@@ 2 Notations, Assumptions and Measures
Notations. The directed communication network is denotes as ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$, where ${\mathcal{N}}=\{1,2,\ldots,n\}$ is the set of nodes, and ${\mathcal{E}}$ is the set of directed edges. An edge $(j,i)\in{\mathcal{E}}$ means that node $i$ can send information to node $i$. We say a directed graph is strongly connected if there is a directed path from any node to all other nodes in the graph. We say that ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$ is the underlying topology of weight matrix $A$ if $A$ is entrywise nonnegative, and $a_{ij}>0$ if and only if $(j,i)\in{\mathcal{E}}$.
Let $\mathds{1}\in\mathbb{R}^{n}$ denote the vector of all-ones. The set $[n]$ denotes the indices $\{1,2,\ldots,n\}$. For a vector $v$, $\mathrm{Diag}(v)$ signifies the diagonal matrix whose diagonal elements are comprised of $v$.
We use matrix $A$ to denote the row-stochastic matrix ($A\mathds{1}=\mathds{1}$) and matrix $B$ to denote the column-stochastic matrix ($B^{\top}\mathds{1}=\mathds{1}$). The Perron vectors [42] for matrices $A$ and $B$, if they exist, are denoted as $\pi_{A}$ and $\pi_{B}$, respectively. We introduce $A_{\infty}:=\mathds{1}\pi_{A}^{\top}$ and $B_{\infty}:=\pi_{B}\mathds{1}^{\top}$. Additionally, we define $n\times d$ matrices
@@
Assumptions. We first specify the underlying graphs (${\mathcal{G}}_{A}$, ${\mathcal{G}}_{B}$) and weight matrices ($A$, $B$) used in stochastic Push-Pull.
@@
Weight matrix $A$ is row-stochastic, with an underlying topology of ${\mathcal{G}}_{A}$. Weight matrix $B$ is column-stochastic, with an underlying topology of ${\mathcal{G}}_{B}$. The traces of $A$ and $B$ are positive.
Assumptions 1 and 2 are commonly used in the analysis of Push-Pull [24, 40, 26]. In our paper, the primary role of these assumptions is to derive Proposition 1. Strictly speaking, replacing these assumptions while retaining Proposition 1 would not affect our main convergence results.
Proposition 1 (Perron Vector, Exponential Decay).
For $A$ and $B$ satisfying Assumption 1 and 2, there exist entrywise positive Perron vectors $\pi_{A},\pi_{B}\in\mathbb{R}^{n}$ such that
@@
Besides, there exists constants $p_{A},p_{B}>0$, $\lambda_{A},\lambda_{B}\in[0,1)$ satisfying
@@
$\displaystyle\mathbb{E}_{k}[\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})]$ $\displaystyle=\nabla f_{i}(x_{i}^{(k)}),\quad\mathbb{E}_{k}[\|\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})-\nabla f_{i}(x_{i}^{(k)})\|_{F}^{2}]\leq\sigma^{2},\quad\forall i\in[n],\,k\geq 0.$
@@
Measures. When Proposition 1 holds, we can define the following constants
$\displaystyle s_{A}:=\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\quad s_{B}:=\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2},\quad c:=n\pi_{A}^{\top}\pi_{B}.$ (2)
In this paper, we use only $s_{A}$, $s_{B}$, $c$, and $n$ to characterize the influence of the mixing matrices. Existing analyses on directed graphs typically employ measures related to the induced matrix norm [25, 38, 39]. In contrast, our measures only require the $\ell_{2}$ norm. An interesting relationship between our measures and other measures is presented in Proposition 2.
@@
Suppose Assumption 1 and 2 hold. We have the following estimate on $s_{A}$, $s_{B}$ and $c$:
$\displaystyle 1\leq s_{A}\leq\frac{1+\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}},$ $\displaystyle\quad 1\leq s_{B}\leq\frac{1+\frac{1}{2}\ln(\kappa_{B})}{1-\beta_{B}},$
$\displaystyle n\max\{\min\pi_{A},\min\pi_{B}\}$ $\displaystyle\leq\,c\leq n\min\{\max\pi_{A},\max\pi_{B}\},$
where $\beta_{A}:=\|\Pi_{A}^{\frac{1}{2}}(A-A_{\infty})\Pi_{A}^{-\frac{1}{2}}\|_{2}\in[0,1)$, $\beta_{B}:=\|\Pi_{B}^{-\frac{1}{2}}(B-B_{\infty})\Pi_{B}^{\frac{1}{2}}\|_{2}\in[0,1),$ $\kappa_{A}:=\max\pi_{A}/\min\pi_{A}$, $\kappa_{B}:=\max\pi_{B}/\min\pi_{B}$ are measures from literature [25, 38, 39], $\Pi_{A}=\mathrm{Diag}(\pi_{A})$ and $\Pi_{B}=\mathrm{Diag}(\pi_{B})$.
@@ 3 Preliminary
In this section, we provide a concise overview of the linear speedup analysis for the Push-Only and Pull-Only algorithms. Furthermore, we discuss the limitations of existing analytical approaches and explain why they cannot be directly extended to the Push-Pull algorithm.
Push-Only and Pull-Only Algorithms Align with Centralized SGD. Most decentralized algorithms, such as decentralized SGD [36], decentralized gradient tracking [43, 14], (sub)gradient-Push [44], Push-Diging [7], (sub)gradient-Pull [9], and FROST [19], are designed to mimic centralized SGD. These algorithms can generally be formulated as:
@@
where $\hat{x}^{(k)}=\sum_{i=1}^{n}v_{i}^{(k)}x_{i}^{(k)}$ is a convex combination of local decision variables, and $\bar{g}^{(k)}=\frac{1}{n}\sum_{i=1}^{n}\nabla F_{i}(x_{i}^{(k)},\xi_{i}^{(k)})$ represents the average gradient. Here, $\epsilon^{(k)}$ denotes an additional error term.
For Push-Only and Pull-Only optimization algorithms, the extra error term is elaborately designed to be zero or to approach zero exponentially fast. Such design ensures that the trajectories of decision variables closely resemble those of centralized SGD. Given the sufficiently small extra error, the linear speedup result naturally follows within the same analytical framework used for centralized SGD. The following is a brief example of how to obtain the formulation (3).
Example 1.
Push-Diging [7]. The algorithm can be written as:
@@
where $V_{k}=\mathrm{Diag}(v_{k})=\mathrm{Diag}(B^{k}\mathds{1})$ is a diagonal matrix and $A_{k}=V_{k+1}^{-1}BV_{k}$. By left-multiplying $n^{-1}v_{k+1}^{\top}$ on both sides of (4a) and omitting some calculation details, we will obtain
@@
By taking $\hat{x}^{(k)}=n^{-1}v_{k}^{\top}{\mathbf{x}}^{(k+1)}$, this is exactly the form of (3).
Figure 1: Distributed logistic regression on synthetic dataset. We use stochastic gradients and a constant learning rate 0.0050.005 0.0050.005. The top three plots illustrate the convergence behavior across different network topologies, while the bottom three plots demonstrate that the size of extra error term rapidly stabilizes and persists as a significant constant. For detailed settings, see 36, Appendix C.
Push-Pull Analysis Suffers from Extra Error Terms. Unlike the aforementioned algorithms, it is generally impossible to reformulate Push-Pull to obtain a centralized SGD-like update. The recursion of Push-Pull is
@@
[…]
$\hat{x}^{(k+1)}=\hat{x}^{(k)}-c\alpha\bar{g}^{(k)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)},$
where $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$ is the projected decision variable, $c=n\pi_{A}^{\top}\pi_{B}$ is a constant. We observe that the error term $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}$ does not vanish but remains a large constant, see Figure 1.
This non-vanishing error term makes achieving linear speedup in Push-Pull highly nontrivial, because in traditional linear speedup analysis, only the noise from $\bar{g}^{(k)}$ is the dominating factor. But here, the error term $\|\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}\|$ introduces additional noise that may dominate the noise in $\bar{g}^{(k)}$.
For certain specific topologies, the noise in the error term can either be ignored or absorbed. For instance, the work in [5] explores a special case known as the $B$-ary tree Push-Pull and demonstrates the linear speedup result on this particular topology. In their setting, $\pi_{A}=\pi_{B}=[1,0,\ldots,0]^{\top}$, which leads to the expression $\pi_{A}^{\top}(I-B_{\infty})=(0,-1,-1,\ldots,-1)$. By setting the loss function at the first node to zero, they obtain the relationship $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}=-n\bar{g}^{(k)}$, which simplifies the process of achieving linear speedup.
Unfortunately, for general topologies, the error term lacks exploitable structure, making direct analysis overly conservative. This leads to a significant overestimation of the noise term, which in turn obstructs the derivation of a linear speedup result. Consequently, the central challenge becomes: How can we accurately characterize the noise within this error term? We tackle this key question in the following section.

### v2

@@ 1.1 Decentralized algorithms over digraphs
The second family consists of algorithms that use only row-stochastic matrices for communication, referred to as the Pull-Only setting. Just as Push-Sum underpins the design of Push-Only algorithms, the Pull-Diag gossip protocol [9] serves as the foundation for Row-Only methods. Building on Pull-Diag, reference [9] adapted the distributed gradient descent algorithm to the Row-Only setting. Subsequently, [18, 19] extended gradient tracking techniques to this framework, while momentum-based Pull-Only gradient tracking methods were developed by [20, 21].
@@ 2 Notations, Assumptions and Measures
Notations. The directed network is denoted by $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}=\{1,2,\ldots,n\}$ is the set of nodes and $\mathcal{E}$ is the set of directed edges. An edge $(j,i)\in\mathcal{E}$ indicates that node $j$ can send information to node $i$. A directed graph is said to be strongly connected if there exists a directed path from any node to every other node in the graph. We say that $\mathcal{G}=(\mathcal{N},\mathcal{E})$ is the underlying topology of the weight matrix $A=[a_{ij}]$ if $A$ is entrywise nonnegative and $a_{ij}>0$ if and only if $(j,i)\in\mathcal{E}$. Let $\mathds{1}\in\mathbb{R}^{n}$ denote the all-ones vector. Set $[n]$ denotes $\{1,2,\ldots,n\}$. For a vector $v$, $\mathrm{Diag}(v)$ denotes the diagonal matrix with $v$ as its diagonal entries.
We use matrix $A$ to denote the row-stochastic matrix (i.e., $A\mathds{1}=\mathds{1}$) and matrix $B$ to denote the column-stochastic matrix (i.e., $B^{\top}\mathds{1}=\mathds{1}$). The Perron vectors [42] of matrices $A$ and $B$, if they exist, are denoted by $\pi_{A}$ and $\pi_{B}$, respectively. We introduce $A_{\infty}:=\mathds{1}\pi_{A}^{\top}$ and $B_{\infty}:=\pi_{B}\mathds{1}^{\top}$. Additionally, we define $n\times d$ matrices
@@
Assumptions. We first specify the underlying graphs ($\mathcal{G}_{A}$, $\mathcal{G}_{B}$) and the corresponding weight matrices ($A$, $B$) used in the stochastic Push-Pull method.
@@
The weight matrix $A$ is row-stochastic with underlying digraph $\mathcal{G}_{A}$, and the weight matrix $B$ is column-stochastic with underlying digraph $\mathcal{G}_{B}$. Both $A$ and $B$ have positive trace.
Assumptions 1 and 2 are standard in the analysis of the Push-Pull method [24, 40, 26]. In our paper, their primary role is to facilitate the derivation of Proposition 1. Strictly speaking, these assumptions could be replaced as long as Proposition 1 remains valid, without affecting the main convergence results.
Proposition 1 (Exponential Decay).
For $A$ and $B$ satisfying Assumptions 1 and 2, there exist entrywise positive Perron vectors $\pi_{A},\pi_{B}\in\mathbb{R}^{n}$ such that
@@
Besides, there exists constants $p_{A},p_{B}>0$, $\lambda_{A},\lambda_{B}\in[0,1)$ such that
@@
$\displaystyle\mathbb{E}_{k}[\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})]$ $\displaystyle=\nabla f_{i}(x_{i}^{(k)})\quad\mbox{and}\quad\mathbb{E}_{k}[\|\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})-\nabla f_{i}(x_{i}^{(k)})\|_{F}^{2}]\leq\sigma^{2}$
for any $i\in[n]$ and $k\geq 0$.
@@
Measures. Various effective metrics have been proposed to capture the impact of digraphs on decentralized stochastic optimization. Let $\Pi_{A}:=\mathrm{Diag}(\pi_{A})$ and $\Pi_{B}:=\mathrm{Diag}(\pi_{B})$. Given a column-stochastic mixing matrix $B$, references [25, 38] introduce $\beta_{B}:=\|\Pi_{B}^{1/2}(B-B_{\infty})\Pi_{B}^{-1/2}\|_{2}\in[0,1)$ as a measure of graph connectivity, and define $\kappa_{B}:=\max\pi_{B}/\min\pi_{B}$ to reflect the deviation of the equilibrium distribution $\pi_{B}$ from the uniform distribution $n^{-1}\mathds{1}$. Similarly, reference [38] introduces $\beta_{A}:=\|\Pi_{A}^{1/2}(A-A_{\infty})\Pi_{A}^{-1/2}\|_{2}\in[0,1)$ and $\kappa_{A}:=\max\pi_{A}/\min\pi_{A}$ to characterize the influence of the row-stochastic mixing matrix $A$ on decentralized algorithms.
However, this paper introduces a new set of metrics to facilitate the convergence analysis of the Push-Pull method. When Proposition 1 holds, we define the following constants:
$\displaystyle s_{A}$ $\displaystyle:=\max\{\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}^{2}\},$ (2)
$\displaystyle s_{B}$ $\displaystyle:=\max\{\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2},\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2}^{2}\},$ (3)
$\displaystyle c$ $\displaystyle:=n\,\pi_{A}^{\top}\pi_{B}.$ (4)
Throughout this paper, we characterize the influence of the mixing matrices using only $s_{A}$, $s_{B}$, $c$, and $n$. A connection between these quantities and the previously discussed metrics is established in Proposition 2.
@@
Suppose Assumption 1 and 2 hold. The following inequalities on $s_{A}$, $s_{B}$ and $c$ hold:
$\displaystyle 1\leq s_{A}\leq\frac{M_{A}^{2}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}},$ $\displaystyle\quad 1\leq s_{B}\leq\frac{M_{B}^{2}(1+\frac{1}{2}\ln(\kappa_{B}))}{1-\beta_{B}},$
$\displaystyle n\max\{\min\pi_{A},\min\pi_{B}\}$ $\displaystyle\leq\,c\leq n\min\{\max\pi_{A},\max\pi_{B}\}.$
where $M_{A}:=\max_{i\geq 0}\|A^{i}-A_{\infty}\|_{2}$, $M_{B}:=\max_{i\geq 0}\|B^{i}-B_{\infty}\|_{2}$.
@@ 3 Preliminary
In this section, we present a concise overview of the linear speedup analysis for the Push-Only and Pull-Only algorithms. We also discuss the limitations of existing analytical approaches and explain why they cannot be directly extended to the Push-Pull algorithm.
3.1 Push-Only and Pull-Only methods align with centralized SGD
Most decentralized algorithms—such as decentralized SGD [36], gradient tracking [31, 32, 43, 44], gradient-Push [45], Push-DiGing [7], gradient-Pull [9], and FROST [19]—are designed to emulate centralized SGD in a decentralized setting. These algorithms can generally be transformed into the following form:
@@ 3.1 Push-Only and Pull-Only methods align with centralized SGD
where $\hat{x}^{(k)}=\sum_{i=1}^{n}v_{i}^{(k)}(x_{i}^{(k)})^{\top}\in\mathbb{R}^{1\times d}$ is a convex combination of the local optimization variables, with $v_{i}^{(k)}$ denoting the combination weights, and $\bar{g}^{(k)}=\frac{1}{n}\mathds{1}^{\top}{\mathbf{g}}^{(k)}$ is the average stochastic gradient. The term $\epsilon^{(k)}$ captures additional error due to the decentralized nature of the algorithm.
For Push-Only and Pull-Only optimization algorithms, the additional error term $\epsilon$ is carefully designed to be either zero or to decay exponentially fast. This design ensures that the trajectories of the local decision variables closely track those of centralized SGD. When the error is sufficiently small, the linear speedup result naturally follows from the same analytical framework used for centralized SGD. Below, we provide a brief example illustrating how to derive the formulation in (5).
Example 1 (Push-Diging [7]).
The Push-Diging algorithm iterates as follows:
@@
where $v_{k}=B^{k}\mathds{1}$, $V_{k}=\mathrm{Diag}(v_{k})$ is a diagonal matrix, $A_{k}=V_{k+1}^{-1}BV_{k}$, and ${\mathbf{g}}^{(k)}$ is the stacked stochastic gradient defined in the notation section. By left-multiplying both sides of (6a) by $n^{-1}v_{k+1}^{\top}$, we obtain
@@
Letting $\hat{x}^{(k)}=n^{-1}v_{k}^{\top}{\mathbf{x}}^{(k)}$, this update exactly matches the form in (5).
Example 2 (FROST [18, 19]).
FROST/PullDiag-GT algorithm only relies on a row-stochastic matrix $A$. Its recursion can be written as follows:
$\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=A{\mathbf{y}}^{(k)}+D_{k+1}^{-1}{\mathbf{g}}^{(k+1)}-D_{k}^{-1}{\mathbf{g}}^{(k)},$ (7b)
where $D_{k}=\mathrm{Diag}(A^{k})$ is a diagonal matrix, ${\mathbf{g}}^{(k)}$ is the stacked stochastic gradient and ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. By left-multiplying both sides of (7a) by $n^{-1}\pi_{A}^{\top}$, we obtain
$n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(k+1)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(k)}-\alpha\bar{g}^{(k)}+\epsilon^{(k)},$
where $\epsilon^{(k)}=\alpha n^{-1}\left(\mathds{1}^{\top}-\pi_{A}^{\top}D_{k}^{-1}\right){\mathbf{g}}^{(k)}$ converges to zero exponentially fast [Lemma 12, [39]]. Letting $\hat{x}^{(k)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(k)}$, this update matches the form  (5).
Figure 1: The extra error does not vanish. The top three plots illustrate the convergence behavior across different network topologies, while the bottom three plots demonstrate that the size of extra error term rapidly stabilizes and persists as a significant constant. The experiment is a logistic regression on synthetic dataset. We use stochastic gradients and a constant learning rate 0.0050.005 0.0050.005, and detailed settings is referred to Appendix C.
3.2 Push-Pull suffers from extra error terms
Unlike the aforementioned algorithms, it is generally impossible to reformulate Push-Pull to obtain a centralized SGD-like update. The recursion of Push-Pull is
@@ 3.2 Push-Pull suffers from extra error terms
where $A$, $B$ are row-stochastic and column-stochastic matrices determined by the topology, $x_{1}^{(0)}=x_{2}^{(0)}=\cdots=x_{n}^{(0)}=x^{(0)}$, ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. Left-multiplying both sides of (8b) by $\mathds{1}^{\top}$, we achieve
$\displaystyle\mathds{1}^{\top}{\mathbf{y}}^{(k+1)}=\mathds{1}^{\top}{\mathbf{y}}^{(k)}+\mathds{1}^{\top}{\mathbf{g}}^{(k+1)}-\mathds{1}^{\top}{\mathbf{g}}^{(k)}=\mathds{1}^{\top}{\mathbf{g}}^{(k+1)},$ (9)
where the last equality holds due to the initialization ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. Next, we left-multiply both sides of (8a) by $\pi_{A}^{\top}$, which yields
$\displaystyle\pi_{A}^{\top}{\mathbf{x}}^{(k+1)}$ $\displaystyle=\pi_{A}^{\top}A{\mathbf{x}}^{(k)}-\alpha\pi_{A}^{\top}B_{\infty}{\mathbf{y}}^{k}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{k}$
$\displaystyle=\pi_{A}^{\top}{\mathbf{x}}^{(k)}-\alpha\pi_{A}^{\top}\pi_{B}\mathds{1}^{\top}{\mathbf{y}}^{k}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{k}.$
Let $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$, and recall from (9) that $\mathds{1}^{\top}{\mathbf{y}}^{(k)}=\mathds{1}^{\top}{\mathbf{g}}^{(k)}$. Then, we have
$\displaystyle\hat{x}^{(k+1)}=\hat{x}^{(k)}-c\alpha\bar{g}^{(k)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)},$ (10)
where $c=n\pi_{A}^{\top}\pi_{B}$ is a constant and $\bar{g}^{(k)}=n^{-1}\mathds{1}^{\top}{\mathbf{g}}^{(k)}$. We observe that the error term $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}$ does not vanish but remains a large constant, see Figure 1.
This non-vanishing error term makes achieving linear speedup in Push-Pull highly nontrivial, because in traditional linear speedup analysis, only the noise from $\|\bar{g}^{(k)}\|^{2}$ is the dominating factor. But here, both the squared error term $\|\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}\|^{2}$ and the cross product $\left\langle\bar{g}^{(k)},\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}\right\rangle$ can introduce additional noise that may dominate the noise in $\bar{g}^{(k)}$.
For certain specific topologies, the error term can be ignored or absorbed. For instance, the work in [5] explores a special case known as the $B$-ary tree Push-Pull (BTPP) and demonstrates the linear speedup result on this particular topology. In BTPP, $\pi_{A}=\pi_{B}=[1,0,\ldots,0]^{\top}$, leading to $\pi_{A}^{\top}(I-B_{\infty})=(0,-1,-1,\ldots,-1)$. By setting the loss function at the first node to zero, $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}=-n\bar{g}^{(k)}$ is achieved, which simplifies the process of achieving linear speedup. Another case is when $A$ is doubly-stochastic; we have $\pi_{A}^{\top}(I-B_{\infty})=0$, which removes the error term.
Unfortunately, for general topologies, the error term lacks exploitable structure. This usually leads to a significant overestimation of the noise term, which in turn obstructs the derivation of a linear speedup result. Therefore, we recognize the central challenge in the analysis of stochastic Push-Pull: How to estimate the noise level in error terms more accurately? We tackle this key question in the following section.